OpenStudy (anonymous):

Find the volume of the solid obtained by revolving the region bounded by y=x^2, y=0, x=0, and x=9 about the y-axis. Can someone explain how to solve this?

OpenStudy (campbell_st):

well the formula is \[V = \pi \int\limits_{f(a)}^{?f(b)} (x^2) dy\] and the graph looks like  so you know x^2 and the question becomes \[V = \pi \int\limits\limits_{0}^{?81} (y) dy\]